# NRGBoost: Energy-Based Generative Boosted Trees

### v1

@@ D.1 Datasets
We use 5 datasets from the UCI Machine Learning Repository (Dheeru & Karra Taniskidou, 2017): Abalone, Physicochemical Properties of Protein Tertiary Structure (from hereon referred to as Protein), Adult, MiniBooNE and Covertype. We also use the California Housing dataset which was downloaded through the Scikit-Learn package Pedregosa et al. (2011) and a downsampled version of the MNIST dataset Deng (2012). Table 7 summarizes the main details of these datasets as well as the approximate number of samples used for train/validation/test for each cross-validation fold.
@@ D.2 DEF and NRGBoost Implementation Details
In our practical implementation of tree based methods we first discretize the input space by binning continuous numerical variables by quantiles. Furthermore we also bin discrete numerical variables in order to keep their cardinalities smaller than 256. This can also be interpreted as establishing a priori a set of discrete values to consider when splitting on each numerical variable and is done for computational efficiency, being inspired by LightGBM (Ke et al., 2017).
@@
For splitting on a categorical variable we once again take inspiration from LightGBM. Rather than relying on one-vs-all splits we found it better to first order the possible categorical values at a leaf according to a pre-defined sorting function and then choose the optimal many-vs-many split as if the variable was numerical. The function used to sort the values is the leaf value function. E.g., for splitting on a categorical variable $x_{i}$ we order each possible categorical value $k$ by $\nicefrac{{\hat{P}(x_{i}=k,X_{-i})}}{{\hat{Q}(x_{i}=k,X_{-i})}}$ where $X_{-i}$ denotes the leaf support over the remaining variables.
@@
We always grow trees in best first order, i.e., we always split the current leaf node that yields the maximum gain in the chosen objective value.
@@
As mentioned in Section 3, we perform a line search to find the optimal step size after each round of boosting in order to maximize the likelihood gain in Equation 45. Because evaluating multiple possible step sizes, $\alpha_{t}$, is inexpensive, we simply do a grid search over 101 different step sizes, split evenly in log space over $[10^{-3},10]$.
@@
Our implementation of the proposed tree-based methods is mostly Python code using the NumPy library (Harris et al., 2020). We implement the tree evaluation and Gibbs sampling in C, making use of the PCG library (O’Neill, 2014) for random number generation. Our code is available in the supplementary material and we will make both the DEF and NRGBoost methods available in an open-source library.
@@ D.3.1 XGBoost
Each model was trained for 1000 boosting rounds on regression and binary classification tasks. For multi-class classification tasks a maximum number of 200 rounds of boosting was used due to the larger size of the datasets and because a separate tree is built at every round for each class. The best model was selected based on the validation set, together with the boosting round where the best performance was attained. The test metrics reported correspond to the performance of the selected model at that boosting round on the test set.
D.3.2 RFDE
We implement the RFDE method (Wen & Hang, 2022) after quantile discretization of the dataset and therefore split at the midpoint of the discretized dimension instead of the original one. When a leaf support has odd cardinality over the splitting dimension a random choice is made over the two possible splitting values. Finally, the original paper does not mention how to split over categorical domains. We therefore choose to randomly split the possible categorical values for a leaf evenly as we found that this yielded slightly better results than a random one vs all split.
For RFDE models we train a total of 1000 trees. The only hyperparameter that we tune is the maximum number of leaves per tree for which we test the values $[2^{6},2^{7},\ldots,2^{14}]$. For the Adult dataset, due to limitations of our tree evaluation implementation we only test values up to $2^{13}$.
D.3.3 ARF
For ARF we used the official python implementation of the algorithm (Blesch & Wright, 2023). We implement the single variable inference for this method ourselves since the python library does not provide density evaluation. Due to memory (and time) concerns we train models with a maximum of 100 trees for up to 10 adversarial iterations. We use a similar grid-search setup to DEF, tuning only the following two parameters:
@@ D.3.3 ARF
  • The minimum number of examples falling on each leaf, for which we test the values $[3,5,10,30,50,100]$. If one value of this parameter doesn’t improve upon the previous, we move on to the next value of the maximum number of leaves (i.e., we exit this loop early).
D.3.4 DEF
We train ensembles with 1000 DET models. Only three hyperparameters are tuned, using three nested loops, each loop running over the possible values of a single parameter in a pre-defined order. These are, in order of outermost to innermost:
  • The maximum number of leaves per tree for which we test the values $[16384,4096,1024,256]$. If the best model for one value of the maximum number of leaves (over the remaining inner-loop parameters) doesn’t improve upon the previous, we stop the hyperparameter tuning. This is done because in most datasets, reducing the number of leaves always leads to a worse model for DEF.
  • The fraction of features to consider when splitting a node. We test the values $[d^{-1/2},d^{-1/4},1]$ with $d$ being the dimension of the dataset. If the best model for one value of this parameter doesn’t improve upon the best model for the previous value, we move on to the next value of the maximum number of leaves (i.e., we exit this loop early).
  • The minimum number of data points that need to be left in each leaf for a split to be considered. We test the values $[0,1,3,10,30]$. If one value of this parameter doesn’t improve upon the previous, we move on to the next value of fraction of features.
D.3.5 NRGBoost
We train NRGBoost models for a maximum of 200 rounds of boosting. We only tune two parameters for NRGBoost Models:
  • The maximum number of leaves for which we try the values $[64,256,1024,4096]$ in order, stopping if performance fails to improve from one value to the next. For the CT dataset we also include 16384 in the values to test.
  • The constant factor by which the optimal step size determined by the line search is shrunk at each round of boosting. This is essentially the "learning rate" parameter. To tune it we perform a Golden-section search for the log of its value using a total of 6 evaluations. The range we use is $[0.01,0.5]$.
For regularization we limit the ratio between empirical data density and model data density on each leaf to a maximum of 2. We noticed that smaller values of this regularization parameter tend to work better but that it otherwise plays a similar role to the shrinkage factor so we opt to tune only the latter.
For sampling we use a sample pool of 80000 samples for all datasets except for Covertype where we use 320000. These values are chosen so that the number of samples in the pool are at minimum similar to the training set size which we found to be a good rule of thumb. At each round of boosting, samples are removed from the pool according to our rejection sampling approach (using Equation 12) after independently removing 10% at random. These samples are replaced by samples from the current model using Gibbs sampling.
The starting point of each NRGBoost model was selected as a mixture model between a uniform distribution (10%) and the product of training marginals (90%) on the discretized input space. We observed that this mixture coefficient does not have much impact on the results however.
D.3.6 TVAE
@@
Table 9: TVAE hyperparameter tuning search space. We set both compress_dims and decompress_dims to have the number of layers specified by num_layers, with hidden_dim hidden units in each layer. We use larger batch sizes and smaller number of epochs for the larger datasets (MBNE, MNIST, CT) since these can take signifcantly longer to run a single epoch.
@@
D.3.7 TabDDPM
We used the official implementation of TabDDPM [^3] https://github.com/yandex-research/tab-ddpm adapted to use our datasets and validation setup. To tune the hyperparameters of TabDDPM we use 50 trials of random search with the same search space that the original authors report in their paper (Kotelnikov et al., 2023).
D.3.8 Forest-Flow
We use the official implementation of Forest-Flow. [^4] https://github.com/SamsungSAILMontreal/ForestDiffusion The only hyperparameters with impact on performance that the Forest-Flow documentation recommends tuning are the number of diffusion timesteps for which a recommended default value of 50 is offered and the number of noise values per sample for which the default value is 100. Due to the fact that Forest-Flow models already take significantly longer to train than the other models, and that that increasing either value would lead to slower training we opted to use these default values.
@@ D.4 Evaluation Setup
Single variable inference
For the single variable inference evaluation, the best models are selected by their discriminative performance on a validation set. The entire setup is repeated five times with different cross-validation folds and with different seeds for all sources of randomness. For the Adult and MNIST datasets the test set is fixed but training and validation splits are still rotated.
@@
For the sampling evaluation we use a single train/validation/test split of the real data (corresponding to the first fold in the previous setup) for training the generative models. The density models used are those previously selected based on their single variable inference performance on the validation set. For TVAE and TabDDPM we directly evaluate their ML Efficiency using the validation data.
@@
For each selected model we sample a train and validation sets with the same number of samples as those used on the original data. For NRGBoost we generate these samples by running 64 chains in parallel with 100 steps of burn in and downsampling their outputs by 30 (for the smaller datasets) or 10 (for MBNE, MNIST and CT). For every synthetic dataset, an XGBoost model is trained using the best hyperparameters found on the real data and using a synthetic validation set to select the best stopping round for XGBoost. The setup is repeated 5 times with different datasets being generated for each method and dataset.
@@
  • To form the synthetic part of the training, validation and test sets for the smaller datasets (AB, CH, PR, AD) we sample data according to the original number of samples in the train, validation and test splits on the real data respectively. Note that this makes the ratio of real to synthetic data 1:4 in the training set. This is deliberate because for these smaller datasets the original validation has few samples and adding extra synthetic data helps the discriminator.
  • For the larger datasets (MBNE, MNIST, CT) we generate the same number of synthetic samples as there are real samples on each split, therefore making every ratio 1:1 because the discriminator is typically already too powerful without adding additional synthetic data.
Because, in contrast to the previous metric, having a lower number of effective samples helps rather than hurts we take extra precautions to not generate correlated data with NRGBoost. We draw each sample by running its own independent chain for 100 steps starting from an independent sample from the initial model which is a rather slow process. The setup is repeated 5 times with 5 different sets of generated samples from each method.
Statistical Distance Measure
We follow a procedure similar to Jolicoeur-Martineau et al. (2024). We min-max scale the numerical variables, and one-hot encode the categorical variables and scale them by 1/2. We use a $L_{1}$ distance in the formulation of the optimal transport problem and use the python POT library to solve it. We sub-sample a maximum of 5000 samples from the original train or test sets and use an equal number of synthetic samples, repeating each experiment 5 times.
@@ D.5 Computational Resources
The experiments were run on a machine equipped with an AMD Ryzen 7 7700X 8 core CPU and 32 GB of RAM. The comparisons with TVAE and TabDDPM further made use of a GeForce RTX 3060 GPU with 12 GB of VRAM.
@@ Appendix E Additional Results
E.1 Computational Effort

### v2

@@ D.1 Datasets
We use 5 datasets from the UCI Machine Learning Repository (Dheeru & Karra Taniskidou, 2017): Abalone, Physicochemical Properties of Protein Tertiary Structure (from hereon referred to as Protein), Adult, MiniBooNE and Covertype. We also use the California Housing dataset which was downloaded through the Scikit-Learn package Pedregosa et al. (2011) and a downsampled version of the MNIST dataset Deng (2012). Table 5 summarizes the main details of these datasets as well as the number of samples used for train/validation/test for the first cross-validation fold.
@@
The toy data distribution used in Figure 2 consists of a mixture of eight isotropic Gaussians with standard deviation of 1, placed around a circle with a radius of 8. To obtain a discrete input domain, we discretize the input space in 100 equally spaced bins between -11 and 11.
@@ D.2 DEF and NRGBoost Implementation Details
In our practical implementation of tree based methods, we first discretize the input space by binning continuous numerical variables by quantiles. Furthermore, we also bin discrete numerical variables in order to keep their cardinalities smaller than 256. This can also be interpreted as establishing a priori a set of discrete values to consider when splitting on each numerical variable and is done for computational efficiency, being inspired by LightGBM (Ke et al., 2017).
@@
For splitting on a categorical variable, we once again take inspiration from LightGBM. Rather than relying on one-vs-all splits, we found it better to first order the possible categorical values at a leaf according to a pre-defined sorting function and then choose the optimal many-vs-many split as if the variable was numerical. The function used to sort the values is the leaf value function. For splitting on a categorical variable, $x_{i}$, we order each possible categorical value $k$ by $\nicefrac{{\hat{P}(x_{i}=k,X_{-i})}}{{\hat{Q}(x_{i}=k,X_{-i})}}$, where $X_{-i}$ denotes the leaf support over the remaining variables.
@@
We always grow trees in best-first order, meaning that we split the current leaf node that yields the maximum gain in the chosen objective value.
@@
As mentioned in Section 3, we perform a line search to find the optimal step size after each round of boosting in order to maximize the likelihood gain in Equation 43. Because evaluating multiple possible step sizes, $\alpha_{t}$, is inexpensive, we simply do a grid search over 101 different step sizes, split evenly in log-space over $[10^{-3},10]$.
@@
Our implementation of the proposed tree-based methods is mostly Python code using the NumPy library (Harris et al., 2020) and Numba. We implement the tree evaluation and Gibbs sampling in C, making use of the PCG library (O’Neill, 2014) for random number generation.
@@ D.3.1 XGBoost
Each model is trained for 1000 boosting rounds on regression and binary classification tasks. For multi-class classification tasks a maximum number of 200 rounds of boosting was used due to the larger size of the datasets and because a separate tree is built at every round for each class. The best model was selected based on the validation set, together with the boosting round where the best performance was attained. The test metrics reported correspond to the performance of the selected model at that boosting round on the test set.
D.3.2 NGBoost
To tune the hyperparameters of NGBoost we use 100 trials of random search with the search space defined in Table 7. We use a maximum of 500 rounds of boosting and early stopping with a patience of 50 rounds. For consistency with the hyperparameter tuning setup for the other models, we select the model that achieves the best $R^{2}$ score in validation.
Table 7: NGBoost hyperparameter tuning search space.
Parameter Distribution
max_depth $\textrm{Uniform}\left([2..10]\right)$
min_samples_leaf $\textrm{Uniform}\left([1..10]\right)$
D.3.3 RFDE
We implement the RFDE method (Wen & Hang, 2022) after quantile discretization of the dataset and therefore split at the midpoint of the discretized dimension instead of the original one. When a leaf support has odd cardinality over the splitting dimension, a random choice is made over the two possible splitting values. Finally, the original paper does not mention how to split over categorical domains. We therefore choose to randomly split the possible categorical values for a leaf evenly as we found that this yielded slightly better results than a random one-vs-all split.
For RFDE models we train a total of 1000 trees. The only hyperparameter that we tune is the maximum number of leaves per tree, for which we test the values $[2^{6},2^{7},\ldots,2^{14}]$. For the Adult dataset, due to limitations of our tree evaluation implementation, we only test values up to $2^{13}$.
D.3.4 DEF
We train ensembles with 1000 DET models. Only three hyperparameters are tuned, using three nested loops, with each loop running over the possible values of a single parameter in a pre-defined order. These are, in order of outermost to innermost:
  1. 1.
    The maximum number of leaves per tree, for which we test the values $[2^{14},2^{12},2^{10},2^{8}]$. For the Adult dataset we use $2^{13}$ instead of $2^{14}$.
  2. 2.
    The fraction of features to consider when splitting a node. We test the values $[d^{-1/2},d^{-1/4},1]$ with $d$ being the dimension of the dataset.
  3. 3.
    The minimum number of data points per leaf. We test the values $[0,1,3,10,30]$.
To speed up the hyperparameter tuning process, we exit each loop early when the best value found for the objective does not improve between consecutive iterations. For example, if for $2^{12}$ leaves per tree, and a feature fraction of $d^{-1/2}$, the objective does not improve when going from 3 minimum points per leaf to 10, we move on to the next value of the feature fraction. Similarly, if the best objective found for a feature fraction of $d^{-1/4}$ (when considering possible values of minimum number of points per leaf) does not improve over that for $d^{-1/2}$, we move on to the next value of number of leaves.
D.3.5 ARF
For ARF we used the official python implementation of the algorithm (Blesch & Wright, 2023). Due to memory (and time) concerns we train models with a maximum of 100 trees for up to 10 adversarial iterations. We use a similar grid-search setup to DEF (with early stopping), tuning only the following two parameters:
  1. 1.
@@ D.3.5 ARF
  2. 2.
    The minimum number of examples per leaf, for which we test the values $[3,5,10,30,50,100]$.
D.3.6 NRGBoost
We train NRGBoost models for a maximum of 200 rounds of boosting. We tune only two parameters, using a similar early stopping setup in the outer loop:
  1. 1.
    The maximum number of leaves, for which we try the values $[2^{6},2^{8},2^{10},2^{12}]$ in order. For the CT dataset we also include $2^{14}$ in the values to test.
  2. 2.
    The constant factor by which the optimal step size determined by the line search is shrunk at each round of boosting. This is essentially the learning rate parameter. To tune it, we perform a Golden-section search for the log of its value using a total of 6 evaluations. The range we use for this search is $[0.01,0.5]$.
For regularization we limit the ratio between empirical data density and model data density on each leaf to a maximum of 2. We observed that smaller values of this regularization parameter tend to perform better, but that it otherwise plays a similar role to the shrinkage factor. Therefore, we choose to tune only the latter.
For sampling, we use a sample pool of 80,000 samples for all datasets except for Covertype where we use 320,000. These values are chosen so that the number of samples in the pool are, at minimum, similar to the training set size, which we found to be a good rule of thumb. At each round of boosting, samples are removed from the pool according to our rejection sampling approach (using Equation 12), after independently removing 10% at random. These samples are replaced by samples from the current model using Gibbs sampling.
The starting point of each NRGBoost model was selected as a mixture model between a uniform distribution (10%) and the product of training marginals (90%) on the discretized input space. However, we observed that this mixture coefficient does not have much impact on the results.
D.3.7 TVAE
@@
Table 8: TVAE hyperparameter tuning search space. We set both compress_dims and decompress_dims to have the number of layers specified by num_layers, with hidden_dim hidden units in each layer. We use larger batch sizes and smaller number of epochs for the larger datasets (MBNE, MNIST, CT) since these can take significantly longer to run a single epoch.
@@
D.3.8 TabDDPM
We used the official implementation of TabDDPM, [^3] https://github.com/yandex-research/tab-ddpm adapted to use our datasets and validation setup. To tune the hyperparameters of TabDDPM we use 50 trials of random search with the same search space used by Kotelnikov et al. (2023).
D.3.9 Forest-Flow
We use the official implementation of Forest-Flow. [^4] https://github.com/SamsungSAILMontreal/ForestDiffusion The only hyperparameters with impact on performance that the Forest-Flow documentation recommends tuning are the number of diffusion timesteps (default of 50) and the number of noise values per sample (default of 100). Due to the fact that Forest-Flow models already take significantly longer to train than the other models, and that increasing either value would lead to slower training, we opted to use these default values.
@@ D.4 Evaluation Setup
Single-variable inference
For the single-variable inference evaluation, the best models are selected by their discriminative performance on a validation set. The entire setup is repeated five times with different cross-validation folds and with different seeds for all sources of randomness. For the Adult and MNIST datasets the test set is fixed but training and validation splits are still rotated.
@@
For the sampling evaluation, we use a single train/validation/test split of the real data (corresponding to the first fold in the previous setup) for training the generative models. The density models used are those previously selected based on their single-variable inference performance on the validation set. For the sampling models (TVAE and TabDDPM) we directly evaluate their ML efficiency using the validation data.
@@
For each selected model we sample a train and validation sets with the same number of samples as those used in the original data. For NRGBoost we generate these samples by running 64 chains in parallel with 100 steps of burn in and downsampling their outputs by 30 (for the smaller datasets) or 10 (for MBNE, MNIST and CT). For every synthetic dataset, an XGBoost model is trained using the best hyperparameters found on the real data and using a synthetic validation set to select the best stopping round for XGBoost. The setup is repeated 5 times with different datasets being generated for each method and dataset.
@@
  • To form the generated part of the training, validation and test sets for the smaller datasets (AB, CH, PR, AD) we sample data according to the original number of samples in the train, validation and test splits on the real data respectivelly. Note that this makes the ratio of real to synthetic data 1:4 in the training set. This is deliberate, because for these smaller datasets the original validation has few samples and adding extra synthetic data helps the discriminator.
  • For the larger datasets (MBNE, MNIST, CT) we generate the same number of synthetic samples as there are real samples on each split, therefore making every ratio 1:1, because the discriminator is typically already too powerful without adding additional synthetic data.
Because, in contrast to the previous metric, having a lower number of effective samples helps rather than hurts, we take extra precautions to not generate correlated data with NRGBoost. We draw each sample by running its own independent chain for 100 steps, starting from an independent sample from the initial model, which is a rather slow process. The setup is repeated 5 times with 5 different sets of generated samples from each method.
@@ D.5 Computational Resources
The experiments were run on a Linux machine equipped with an AMD Ryzen 7 7700X 8 core CPU and 32 GB of RAM. The comparisons with TVAE and TabDDPM additionally made use of a GeForce RTX 3060 GPU with 12 GB of VRAM.
@@ Appendix E Additional Results
E.1 Statistical Distance Measure
As an additional measure of the statistical dissimilarity between data and model distributions, we evaluate a Wasserstein distance using a similar setup to Jolicoeur-Martineau et al. (2024). Namely, we min-max scale numerical variables, and one-hot encode categorical variables, scaling the latter by 1/2. We use a $L_{1}$ distance in the formulation of the optimal transport problem, which we solve using the POT python library. Since finding the optimal solution scales cubically with the sample size in the worst case, we sub-sample at most 5,000 samples from the original train or test sets and use an equal number of synthetic samples. We repeat the evaluation for each method and dataset 5 times, using different synthetic data and also different subsampling seeds for the real data (where applicable). Our results for the distance between empirical training and test distribution and synthetic samples from each model are reported in Tables 9 and 10 respectively.
AB CH PR AD MBNE MNIST CT
AB CH PR AD MBNE MNIST CT
We found these results to be somewhat sensitive to the choice of normalization of the numericals and categoricals, as well as to the choice of distance in this normalized space. Despite NRGBoost not being as dominant as in the discriminator measure, it still fares well, obtaining the best average rank over all datasets and experiments.
E.2 Uncertainty Estimation in Regression Tasks
One advantage of density models over discriminative regression models like XGBoost is that they estimate the full conditional distribution, $p(y|\mathbf{x})$, rather than provide a point estimate. In Section 5.1, we focused on a metric that evaluates the point estimate $E[y|\mathbf{x}]$. To further evaluate the quality of the models’ probabilistic predictions, in Table 11 we report two other metrics:
  • MAE: We compute the average Mean Absolute Error given a point estimate of $y$. For NGBoost and NRGBoost, we use the median of the estimated $q(y|\mathbf{x})$ since this is the point estimate that minimizes MAE. Note that, while it is possible to train XGBoost to directly optimize this metric instead of MSE, models estimating $q(y|\mathbf{x})$ are flexible to choose a different point estimate post-training.
  • (C)RPS: The Continuous Ranked Probability Score is a proper scoring rule commonly used to evaluate a probabilistic forecast (Gneiting & Raftery, 2007). For both NGBoost and NRGBoost, we evaluate this score analytically for the forecasted $q(y|\mathbf{x}_{i})$ of every test point. Because the Abalone dataset has discrete numerical targets, we compute RPS for this dataset instead. For this purpose, the continuous distribution obtained with NGBoost is first discretized by rounding to the nearest discrete value within the domain of the target.
Note that we omit the results for ARF from this analysis because the Python implementation does not provide neither conditional nor unconditional density evaluation at the time of writing. For the results reported in Table 1, we implemented our own evaluation of $E[y|\mathbf{x}]$ for these models. Extending this implementation to compute the median or the CRPS analytically is, however, not trivial due to ARF fitting more complex continuous distributions at the leaves than the other tree-based models.
On MAE, we find that NRGBoost outperforms XGBoost convincingly in the California Housing and Protein datasets, in part due to the use of a better point estimate for this metric. The Abalone dataset is the only dataset where predicting the mean of the forecasting distribution would outperform the median for MAE. We also find that NRGBoost achieves a lower (C)RPS than NGBoost on all datasets, which we attribute to the parametric choice made for NGBoost not necessarily being the most well-suited for these datasets. This highlights a key drawback of parametric approaches compared to non-parametric ones: they require domain expertise to avoid suboptimal results.
Table 11: (C)RPS and MAE results. The reported values are means and standard errors over 5 cross-validation folds. The best method for each dataset is highlighted in bold and other methods that are not significantly worse (as determined by a paired $t$-test at a 95% confidence level) are underlined.
AB CH PR
RPS $\downarrow$ MAE $\downarrow$ CRPS $\downarrow$ MAE $\downarrow$ CRPS $\downarrow$ MAE $\downarrow$
XGBoost N/A 1.524 ±0.032 N/A 0.290 ±0.009 N/A 2.282 ±0.045
NGBoost 1.094 ±0.020 1.525 ±0.032 0.238 ±0.013 0.314 ±0.011 1.976 ±0.018 2.665 ±0.057
NRGBoost 1.075 ±0.011 1.569 ±0.025 0.201 ±0.008 0.276 ±0.010 1.490 ±0.023 2.063 ±0.036
E.3 Inference with Missing Data
Generative density models can be used more flexibly for inference than their discriminative counterparts. In this section we present a case study on a problem of inference in the presence of a missing covariate that highlights this advantage.
For each type of task, we choose a dataset and remove an important feature from the respective test set (Longitude for the California dataset, education for the Adult dataset and Elevation for the Covertype dataset). We then attempt to predict the value of the target variable using XGBoost with the help of several imputation techniques that are popular with data science practitioners:
  • Mean and median imputation (for a missing numerical feature).
  • Mode imputation (for a missing categorical feature).
  • Imputation with the mean or mode of the 5 nearest neighbors on the training data. Note that this imputation requires one to have access to the entire training set at test time.
NRGBoost can handle this type of problem in a principled manner by marginalizing over the possible values of the missing feature, rather than relying on ad-hoc imputation. Our results, in Table 12, show that this approach is able to outperform imputation, even on datasets where a performance gap between NRGBoost and XGBoost existed on the full data.
Table 12: Results for inference with a missing covariate. The original results of each method when having access to the full data are also reported for reference. The best approach for dealing with missing data in each scenario is highlighted in bold.
Model Imputation CH ($R^{2}$uparrow) AD (AUC $\uparrow$) CT (Accuracy $\uparrow$)
XGBoost Full Data 0.849 ±0.009 0.927 ±0.000 0.971 ±0.001
Mean -0.283 ±0.107 N/A 0.610 ±0.004
Median/Mode -0.117 ±0.107 0.914 ±0.003 0.621 ±0.002
KNN (K=5) 0.150 ±0.107 0.910 ±0.003 0.883 ±0.001
NRGBoost Full Data 0.850 ±0.011 0.920 ±0.001 0.948 ±0.001
Marginalization 0.773 ±0.010 0.920 ±0.001 0.923 ±0.001
E.4 Other Challenging Tasks
In order to test the limits of NRGBoost, we compare its performance to XGBoost on the following two challenging tasks taken from Gorishniy et al. (2021):
  • ALOI: This is an image classification dataset with 128 features and 100k examples distributed evenly over 1000 classes. Because our implementation of NRGBoost is limited to variables with a cardinality of 255, we consider only data for the first 250 classes. This still constitutes 25 times more classes than MNIST, presenting a challenge to both discriminative and generative models.
  • Microsoft: This regression task contains both a large number of samples (1.2M) and features (136), making it one order of magnitude larger than MNIST and Covertype in terms of the product of the two. Because of this, we found that our usual hyperparameter tuning protocol for NRGBoost to be impractical and had to resort to manual tuning instead. The best model we found had 256 leaves and was trained for 500 boosting iterations with a maximum ratio in each leaf of 4 and a shrinkage factor of 0.1.
We use the same train, validation and test splits used in Gorishniy et al. (2021) and report the same metrics in Table 13. We find that NRGBoost outperforms XGBoost in the ALOI dataset but that there is still a significant gap between the best NRGBoost model that we were able to find on the Microsoft dataset and XGBoost.
Table 13: Comparison of XGBoost and NRGBoost on ALOI and Microsoft datasets.
ALOI (Accuracy $\uparrow$) Microsoft (RMSE $\downarrow$)
XGBoost 0.943 0.744
NRGBoost 0.962 0.787
E.5 MCMC Chain Convergence on MNIST
Figure 5 shows the convergence of Gibbs sampling from an NRGBoost model. In only a few samples each chain appears to have converged to the data manifold after starting at a random sample from the initial model (a mixture between the product of training marginals and a uniform). Note how consecutive samples are autocorrelated. In particular it can be rare for a chain to switch between two different modes of the distribution (e.g., switching digits) even though a few such transitions can be observed.
Figure 5: Downsampled MNIST samples generated by Gibbs sampling from an NRGBoost model. Each row corresponds to an independent chain initialized with a sample from the initial model $f_{0}$ (first column). Each column represents a consecutive sample from the chain.
E.6 Computational Effort
E.6.1 Training